caos lock shit T Multiloop Circuits 2010 Make a sketch showing the junctions and loops for the analysis of an unbalanced Wheatstone bridge (see your textbook). How many simultaneous equations would be needed for analysis? 2. 3. Two batteries with emfs of 1.5 V and 1.2 V and internal resistances of 0.5 12 and 0.8 Q, respectively, are connected in parallel. (a) Determine the current they would deliver to an external resistance of 2 2. (b) Would more current be delivered if only one battery were used? If so, which one? Explain
Answer #1

3)

Given,

V1 = 1.5 V ; V2 = 1.2 V ; r1 = 0.5 Ohm ; r2 = 0.8 Ohm

2( i1 + i2) + 0.8 i2 = 1.2

2i1 + 2.8i2 = 1.2

2 (i1 + i2) + 0.5 i1 = 1.5

2.5 i1 + 2 i2 = 1.5

solving the above eqn we get

i1 = (1.5 - 2i2)/2.5

2/2.5 (1.5 - 2i2) + 2.8i2 = 1.2

i1 = 1.16 A and i2 = -0.4 A

So the current through 2 Ohm is:

I = 1.16 - 0.4 = 0.76 A

Hence, I = 0.76 A

b)No current will be delivered if we use only battery in the circuit